# Enhancing Domain-Specific Encoder Models with LLM-Generated Data: How to Leverage Ontologies, and How to Do Without Them

### v1

@@ Abstract
We investigate the use of LLM-generated data for continual pretraining of encoder models in specialized domains with limited training data, using the scientific domain of invasion biology as a case study. To this end, we leverage domain-specific ontologies by enriching them with LLM-generated data and pretraining the encoder model as an ontology-informed embedding model for concept definitions. To evaluate the effectiveness of this method, we compile a benchmark specifically designed for assessing model performance in invasion biology. After demonstrating substantial improvements over standard LLM pretraining, we investigate the feasibility of applying the proposed approach to domains without comprehensive ontologies by substituting ontological concepts with concepts automatically extracted from a small corpus of scientific abstracts and establishing relationships between concepts through distributional statistics. Our results demonstrate that this automated approach achieves comparable performance using only a small set of scientific abstracts, resulting in a fully automated pipeline for enhancing domain-specific understanding of small encoder models that is especially suited for application in low-resource settings and achieves performance comparable to masked language modeling pretraining on much larger datasets.
@@ 1 Introduction
Transformer encoder models such as BERT Devlin et al. (2019) and its successors (e.g., Liu et al., 2019, He et al., 2021a, Warner et al., 2024) have consistently achieved state-of-the-art results across various text-based tasks, mainly enabled by pretraining with masked language modeling (MLM) or replaced token detection Clark et al. (2020) on large-scale general-domain corpora, such as Wikipedia and BookCorpus Zhu et al. (2015).
@@
While transformer encoders offer an optimal balance between performance and efficiency, their full effectiveness in specialized domains, such as scientific text processing, is often enabled by additional pretraining on domain-specific corpora Beltagy et al. (2019); Jeong and Kim (2022), proven highly effective in fields where extensive domain-specific data is available (e.g., biomedical text processing Gu et al., 2021). However, in more specialized disciplines with limited training data, the potential of this approach diminishes, highlighting the need for alternative methods of injecting domain knowledge during pretraining.
To this end, we explore the use of domain-specific ontologies for continual pretraining of encoder models in the scientific domain of invasion biology. Recognizing that ontologies may not always be available, we also investigate the extent to which their knowledge can be replaced by LLM-extracted information derived from scientific abstracts.
Our focus on ontologies is grounded in the fact that they exist in many fields with otherwise low data availability, while at the same time containing precise, domain-specific and structured knowledge curated by domain experts (e.g., Walls et al., 2014; Girón et al., 2023; Algergawy et al., 2025), thus making them a valuable resource for pretraining.
Our proposed approach enriches these ontologies with LLM-generated data in the form of concept definitions, followed by pretraining the encoder model as an ontology-informed definition embedding model. Specifically, we employ a triplet margin loss that enforces definitions of either the same or similar concepts to be placed at nearby positions in the embedding space, thus enabling the model to develop a structured understanding of domain-specific entities and their interconnections.
Having demonstrated the effectiveness of our approach, we further explore its applicability in domains where no ontology is available. To this end, we develop a pipeline that automatically extracts relevant concepts from scientific abstracts using an LLM, generates domain-specific definitions, and identifies relationships between concepts based on distributional statistics. Our results indicate that these automatically generated substitutes for ontological components can achieve performance comparable to ontology-based pretraining.
By compiling a benchmark for evaluating models in the invasion biology domain - consisting of four tasks from three existing studies Brinner et al. (2022, 2024); Brinner and Zarrieß (2025) - we demonstrate that our proposed pretraining approach achieves performance comparable to traditional masked language modeling (MLM) on scientific abstracts, thus establishing our method as a viable alternative. Furthermore, we show that our approach, which is applied solely to the model’s CLS token output, can be combined seamlessly with MLM pretraining, leading to significantly improved performance compared to using each method individually, thus indicating a complementary effect on the model’s domain understanding.
The resulting method requires only 5,000 scientific abstracts in combination with an ontology, or 15,000 abstracts independently to match the performance of models pretrained on 14 million abstracts in the broader biomedical domain Gu et al. (2021), underscoring its potential as an effective solution for low-resource settings.
@@ 2 Related Work
Continual Pretraining.
Continual pretraining is an effective and efficient approach to make LLMs robust against new, ever-changing data that differs from its original pretraining Wu et al. (2024); Zhou et al. (2024); Parmar et al. (2024); Shi et al. (2024), that can also further enhance an LLM’s domain specific effectiveness Gururangan et al. (2020); Gong et al. (2022); Xie et al. (2023); Çağatay Yıldız et al. (2025) and specializes in "improving knowledge transfer to downstream tasks" Wang et al. (2024), such as scientific text processing in this study. The phenomenon of catastrophic forgetting poses significant risk in continual pretraining Li and Lee (2024); Ibrahim et al. (2024); Cossu et al. (2024), wherein a continually learned model forgets knowledge from previous training. In this study we perform continual pretraining to specialize a model in the narrow domain of invasion biology and discuss the risk of catastrophic forgetting in Section 6.
Using Ontological Knowledge.
@@
Ontologies and knowledge graphs (KGs) provide a structured representation of domain knowledge in the form of unique entities and precise relations between them, contrasting the distributed and often less precise knowledge representation within neural networks. To bridge this gap, various methods have been proposed to integrate structured knowledge into transformer models. While some approaches incorporate KG information during inference Zhang et al. (2019); Peters et al. (2019); He et al. (2020), the majority of approaches focus on creating KG-informed pretraining methods, for example by performing MLM pretraining that incorporates knowledge about entities Shen et al. (2020); Zhang et al. (2021), performing MLM pretraining on sentences derived from KG triples Lauscher et al. (2020); Moiseev et al. (2022); Liu et al. (2022); Sahil and Kumar (2023); Omeliyanenko et al. (2024), designing auxiliary classification tasks based on ontological knowledge Wang et al. (2021a); Glauer et al. (2023) or by creating contrastive ontology-informed sentence embedding methods Wang et al. (2021b); Ronzano and Nanavati (2024). Our approach aligns most closely with the latter but extends it into a broader framework that incorporates not only relationships between concepts but also LLM-derived knowledge about individual concepts, even in the absence of explicit relations, thus creating a more informative and flexible pretraining process.
Using LLM-Generated Data.
@@
Using LLM-generated data is an appealing approach for model pretraining and/or fine-tuning Long et al. (2024), especially in specialized domains with little available training data. Many studies explore the potential of LLM-generated or LLM-annotated data to enhance task-specific performance, both for encoder models Kruschwitz and Schmidhuber (2024); Kuo et al. (2024); Wagner et al. (2024) and decoder architectures Ren et al. (2024); Lee et al. (2024).
@@
Beyond task-specific fine-tuning, synthetic data has also been investigated for task-agnostic pretraining. While this approach has shown promise for general-domain models Alcoba Inciarte et al. (2024); Yang et al. (2024); McKinzie et al. (2025), its application in domain-specific pretraining remains relatively underexplored (e.g., Yuan et al., 2024).
@@
Despite its advantages, synthetic data introduces risks, including potential performance degradation compared to human-generated data - a phenomenon known as model collapse Shumailov et al. (2024), prompting studies aimed at mitigating this effect, especially for autoregressive LLMs Bertrand et al. (2024); Gerstgrasser et al. (2024); Zhang et al. (2024); Zhu et al. (2024). We discuss the differences between pretraining on scraped vs generated data in Section 6.
@@ 3.1 Similarity-Based Pretraining
We propose a contrastive triplet margin loss for continual pretraining of an encoder model, refining it as an embedding model for concept definitions by teaching it to place definitions of the same concept or definitions of related concepts to similar positions in the embedding space, thus enabling the model to capture both the meaning and distinctions between domain-specific concepts effectively.
Our method operates on a dataset of domain-relevant concepts $\mathcal{C}=\{C_{1},C_{2},...\}$, each in combination with multiple natural language concept definitions $\mathcal{D}=\{(d_{1,1},d_{1,2},...),(d_{2,1},d_{2,2},...),...\}$. Also, we optionally incorporate a set of tuples indicating pairs of related concepts =$\mathcal{R}=\{(C_{i},C_{j}),...\}$ to increase the model’s domain understanding beyond knowledge of individual entities.
The core training scheme is as follows: Given two concepts $C_{i}$ and $C_{j}$ from the dataset, we train the model to embed two definitions of concept $C_{i}$ to nearby locations in the embedding space while positioning a definition of $C_{j}$ further away, thereby teaching the model to understand and differentiate between dissimilar concepts. This is achieved by sampling two definitions $d_{i,1}$ and $d_{i,2}$ that define concept $C_{i}$, and one definition $d_{j,1}$ that defines concept $C_{j}$. These definitions are then mapped into high-dimensional embeddings using our model $M$:
$e_{i,1}=M(d_{i,1})$
$e_{i,2}=M(d_{i,2})$
$e_{j,1}=M(d_{j,1})$
@@
$L=\textrm{relu}(||e_{i,1}-e_{i,2}||-||e_{i,1}-e_{j,1}||+1)$
In this contrastive loss formulation, $d_{i,1}$ serves as an anchor, with $d_{i,2}$ being the positive and $d_{j,1}$ being the negative with respect to that anchor. The loss function thus penalizes cases in which the distance between the anchor and the positive (i.e., two definitions defining the same concept) is not at least one unit (a margin hyperparameter) smaller than the distance between the anchor and the negative.
Rather than explicitly sampling individual triplets (anchor, positive, and negative), we optimize the loss computation by leveraging in-batch negatives. Specifically, for a batch of $n$ concepts, we sample only two definitions - an anchor and a positive - for each concept. The definitions of the remaining concepts in the batch then serve as negative samples. This strategy significantly increases the number of triplets contributing to the loss: for each anchor-positive pair, $2\cdot(n-1)$ triplets are generated by pairing the anchor with all possible negatives, which can be further doubled to $4\cdot(n-1)$ by swapping the roles of the anchor and positive sample. This substantial increase in triplets enhances model performance, as the loss function quickly reaches zero for many triplets after just a few epochs due to the model’s rapidly improving embedding capabilities. Consequently, the larger number of triplets increases the likelihood of encountering more informative gradient signals, ultimately leading to more effective embeddings.
@@ 3.2 Concept Relatedness
The current loss formulation encourages the model to map similar definitions (i.e., those defining the same concept) to nearby positions in the embedding space. While this enhances the model’s ability to differentiate between concepts, a deeper understanding of the domain also requires learning relationships between different concepts. Therefore, we extend our loss formulation by incorporating additional triplets that capture concept relatedness.
Specifically, if two concepts $C_{i}$ and $C_{j}$ are in the same batch and $(C_{i},C_{j})\in\mathcal{R}$, we treat their definitions as additional positive pairs within the loss function, while definitions of all unrelated concepts serve as negatives. This implicitly introduces a ranking effect, since definitions of related concepts still function as negatives for the definition that defines the same concept, ensuring that these are embedded more closely together than definitions of related concepts. Simultaneously, related concepts are encouraged to be positioned closer in the embedding space than unrelated concepts.
@@ 3.3 Pretraining Loss Combination
Our proposed pretraining loss is applied to the CLS token representation, allowing seamless integration with other pretraining losses that target the remaining token embeddings like masked language modeling. This is especially interesting in light of recent models being trained exclusively with MLM loss Warner et al. (2024), since the traditional next sentence prediction loss from BERT did not lead to significant performance gains Liu et al. (2019). Consequently, our method presents a more sophisticated approach for infusing knowledge into the CLS token representation, offering a potentially enhanced downstream task performance when used.
@@ 4 Ontology-Informed Pretraining
This section details the application and evaluation of our proposed method, using domain-specific ontologies for dataset creation. Our experiments focus on the scientific domain of invasion biology, a specialized subfield of biodiversity research that investigates non-native species, their introduction pathways, ecological impacts, and management strategies to mitigate their effects on ecosystems Jeschke and Heger (2018).
@@ 4.1 Dataset Creation
Our approach involves constructing a domain-specific dataset consisting of concepts, definitions and concept relations in the target domain. To this end, we use two ontologies that address the target domain: the INBIO ontology Algergawy et al. (2025), which captures concepts relevant to invasion biology, and the ENVO ontology Buttigieg et al. (2013, 2016), which provides a structured representation of environmental and ecological concepts.
From these ontologies, we extract concept-definition pairs for all concepts that have a corresponding definition along with relational links. Additionally, we use a large language model, LLaMA-3-8B-Instruct Grattafiori et al. (2024), to generate five additional definitions per concept. The original ontology definition serves as context during generation to ensure that the new definitions accurately reflect the domain-specific meaning.
We compare our proposed pretraining approach to traditional MLM pretraining on sentences extracted from scientific abstracts. We leverage an existing index of paper titles in the field of invasion biology Mietchen et al. (2024) and employ a web scraper to retrieve their abstracts, resulting in a final collection of 37,786 paper titles and abstracts.
@@ 4.2 Model Pretraining
In our experiments, we perform continual pretraining on a DeBERTa-base model He et al. (2021b) by leveraging three different pretraining strategies:
@@
    Masked language modeling (MLM) pretraining: We evaluate the effectiveness of traditional MLM pretraining with a masking probability of 0.25, applied to either abstract sentences, generated definitions, or a combined dataset of both.
@@
    Simmilaity (SIM) pretraining: As described in Section 3, we pretrain the model using our proposed similarity-based approach, leveraging the extracted and generated definitions for the ontology concepts.
@@
    Combined pretraining: To investigate potential synergies between MLM and SIM pretraining, we apply both strategies concurrently by performing two backward passes - one for each loss function - for each parameter update.

### v2

@@ Abstract
We investigate the use of LLM-generated data for continual pretraining of transformer encoder models in specialized domains with limited training data, using the scientific domain of invasion biology as a case study. To this end, we leverage domain-specific ontologies by enriching them with LLM-generated data and pretraining the encoder model as an ontology-informed embedding model for concept definitions. To evaluate the effectiveness of this method, we compile a benchmark specifically designed for assessing model performance in invasion biology. After demonstrating substantial improvements over standard MLM pretraining, we investigate the feasibility of applying the proposed approach to domains without comprehensive ontologies by substituting ontological concepts with concepts automatically extracted from a small corpus of scientific abstracts and establishing relationships between concepts through distributional statistics. Our results demonstrate that this automated approach achieves comparable performance using only a small set of scientific abstracts, resulting in a fully automated pipeline for enhancing domain-specific understanding of small encoder models that is especially suited for application in low-resource settings and achieves performance comparable to masked language modeling pretraining on much larger datasets.
@@ 1 Introduction
Transformer encoder models such as BERT devlin2019bert and its successors (e.g., liu2019roberta, he2021debertav3, warner2024modernbert) have consistently achieved state-of-the-art results across a wide range of NLP tasks. These successes are largely driven by large-scale pretraining on general-domain corpora such as Wikipedia and BookCorpus zhu2015books, using objectives like masked language modeling (MLM) or replaced token detection clark2020electra.
@@
While transformer encoders offer an optimal balance between performance and efficiency, their full effectiveness in specialized domains - such as scientific text processing - is often enabled by additional pretraining on domain-specific corpora beltagy2019scibert; 9791256, proven highly effective in fields where extensive domain-specific data is available (e.g., biomedical text processing Gu_2021). However, in specialized disciplines with limited resources, the potential of this approach diminishes, highlighting the need for alternative methods of injecting domain knowledge during pretraining.
To address this challenge, we propose a novel method for continual pretraining of transformer encoder models that leverages a set of domain-relevant concepts and their corresponding definitions as the core training resource. These concepts can be drawn from domain-specific ontologies (i.e., data structures containing precise, domain-specific and structured knowledge curated by domain experts walls2014semantics; 10.1093/sysbio/syad025; alsayed_algergawy_2025_14864244) or extracted from texts using LLMs. Using this resource, we pretrain the model as an embedding model for concept definitions, encouraging definitions of identical or related concepts to occupy nearby positions in the embedding space and thus enabling the model to develop a structured understanding of domain-specific entities and their interconnections.
In this process, we perform an extensive analysis of the effectiveness of incorporating different types of information, moving from using domain-relevant concepts extracted from the ontologies towards a fully unsupervised pipeline with LLM-extracted concepts.
Contributions:
1) We validate the effectiveness of our embedding-based pretraining approach using ontology-derived concepts and LLM-generated definitions, establishing it as a viable alternative to traditional MLM pretraining.
2) We identify the benefits of incorporating concept relatedness by integrating ontological relationship links into the pretraining objective.
3) We explore the possibility of combining our pretraining strategy with traditional MLM pretraining, demonstrating strong synergistic effects that vastly improve downstream performance.
4) We create and evaluate a fully unsupervised pipeline by replacing ontology-derived concepts with LLM-extracted concepts from scientific abstracts. By also using distributional statistics as a concept relatedness indicator, we remove the dependency on manually curated ontologies.
5) We analyze performance of our unsupervised approach across varying dataset sizes, showing that it consistently outperforms MLM pretraining, even when trained on significantly less data.
6) We focus on the use of synthetic data in the form of LLM-generated concept definitions and analyze model collapse shumailov2024curserecursiontraininggenerated, demonstrating that our proposed pretraining scheme is much less susceptible to this issue compared to classical mask language modeling.
Due to the extensiveness of our multi-step experiments and analysis, we focus on a single, representative domain: invasion biology. This field exemplifies a complex, specialized area of scientific research with limited unsupervised pretraining data or annotated resources. To support evaluation, we compile a new benchmark from three existing studies brinner2022linking; brinn2024weaklyclaim; brinner2025efficient, covering a diverse set of tasks that collectively provide a comprehensive assessment of model performance in this domain.
@@ 2 Related Work
Continual Pretraining is an effective and efficient approach to make LMs robust against new, ever-changing data that differs from its original pretraining wu2024continuallearninglargelanguage; zhou2024continuallearningpretrainedmodels; parmar2024reusedontretrainrecipe; shi2024continuallearninglargelanguage, enhances an LLM’s domain specific effectiveness gururangan2020dontstoppretrainingadapt; gong-etal-2022-continual; xie2023efficientcontinualpretrainingbuilding; yıldız2025investigatingcontinualpretraininglarge and improves knowledge transfer to downstream tasks 10444954.
@@
Ontologies and Knowledge Graphs (KGs) have been explored as resource for continual pretraining since they provide a structured representation of domain knowledge in the form of unique entities and precise relations between them, contrasting the distributed and often less precise knowledge representation within neural networks. To bridge this gap, various methods have been proposed to integrate structured knowledge into transformer models. While some approaches incorporate KG information during inference zhang2019ernie; peters2019knowledge; he2020bert, the majority of approaches focus on creating KG-informed pretraining methods, for example by performing MLM pretraining that incorporates knowledge about entities shen2020exploiting; zhang2021ebert, performing MLM pretraining on sentences derived from KG triples lauscher2020common; moiseev2022skill; liu2022enhancing; sahil2023leveraging; Omeliyanenko2024preadapter, designing auxiliary classification tasks based on ontological knowledge wang2021k; glauer2023ontology or by creating contrastive ontology-informed sentence embedding methods wang2021kepler; ronzano2024ontology. Our approach aligns most closely with the latter but extends it into a broader framework that incorporates not only relationships between concepts but also LLM-derived knowledge about individual concepts by incorporating synthetic concept definitions, thus creating a more informative and flexible pretraining process that is not reliant on the presence of ontologies.
@@
Using Synthetic Data for model pretraining and/or fine-tuning is an appealing prospect long2024llmsdrivensyntheticdatageneration, especially in specialized domains with little available training data. Many studies explore the potential of LLM-generated or LLM-annotated data to enhance task-specific performance, both for encoder kruschwitz-schmidhuber-2024-llm; kuo2024llmgenerateddataequalrethinking; wagner2024powerllmgeneratedsyntheticdata and decoder architectures ren2024learn; lee2024llm2llmboostingllmsnovel.
@@
Beyond task-specific fine-tuning, synthetic data has also been investigated for task-agnostic pretraining. While this approach has shown promise for general-domain models alcoba2024utility; yang2024syntheticcontinuedpretraining; 10.1007/978-3-031-73397-0_18, its application in domain-specific pretraining remains relatively underexplored (e.g., yuan2024continuedpretrainedllmapproach).
@@
Despite its advantages, synthetic data introduces risks, including potential performance degradation compared to human-generated data - a phenomenon known as model collapse shumailov2024curserecursiontraininggenerated, prompting studies aimed at mitigating this effect, especially for autoregressive LLMs bertrand2024stabilityiterativeretraininggenerative; gerstgrasser2024modelcollapseinevitablebreaking; zhang2024regurgitativetrainingvaluereal; zhu2024synthesizetextdatamodel. In Section 6, this phenomenon will be further discussed in the context of our own experiments.
@@ 3.1 Similarity-Based Pretraining
We propose a pretraining strategy for transformer encoder models, training them as embedding models for concept definitions by teaching it to place definitions of the same concept or definitions of related concepts to similar positions in the embedding space, thus enabling the model to capture both the meaning and distinctions between domain-specific concepts effectively. A comparable strategy has previously proven effective for training specialized embedding models on scientific abstracts, where it substantially improved semantic encoding capabilities brinner2025semcse, thus suggesting that a related approach may provide an effective means of enforcing semantic understanding of relevant domain knowledge.
[…]
The core training scheme is as follows: Given two concepts $C_{i}$ and $C_{j}$ from the dataset, we train the model to embed two definitions of concept $C_{i}$ to nearby locations in the embedding space while positioning a definition of $C_{j}$ further away, thereby teaching the model an understanding of the different concepts. This is achieved by sampling two definitions $d_{i,1}$ and $d_{i,2}$ that define concept $C_{i}$, and one definition $d_{j,1}$ that defines concept $C_{j}$. These definitions are then mapped into the high-dimensional embedding space using the model $M$, resulting in embeddings $e_{i,1}$, $e_{i,2}$ and $e_{j,1}$.
@@
[…]
In this triplet loss formulation, $d_{i,1}$ serves as an anchor, with $d_{i,2}$ being the positive and $d_{j,1}$ being the negative with respect to that anchor. The loss function thus penalizes cases in which the distance between the anchor and the positive (i.e., two definitions defining the same concept) is not at least one unit (a margin hyperparameter) smaller than the distance between the anchor and the negative.
Rather than explicitly sampling individual triplets (anchor, positive, and negative), we optimize the loss computation by leveraging in-batch negatives, thus only sampling an anchor and a positive for each concept and using all definitions from other concepts within the batch as negatives. This strategy - in combination with switching the roles of anchor and positive - significantly increases the number of triplets contributing to the loss, leading to $4\cdot(n-1)$ triples per anchor-positive pair with a batch-size of $n$. This substantial increase in triplets enhances model performance, as the loss function quickly reaches zero for many triplets after just a few epochs due to the model’s rapidly improving embedding capabilities. Consequently, the larger number of triplets increases the likelihood of encountering more informative gradient signals, ultimately leading to more effective embeddings.
@@ 3.2 Concept Relatedness
The current loss formulation encourages the model to map similar definitions (i.e., those defining the same concept) to nearby positions in the embedding space. While this enhances the model’s ability to differentiate between concepts, a deeper understanding of the domain also requires learning relationships between different concepts, which might otherwise be learned only implicitly through the similarity between their definitions. Therefore, we extend our loss formulation by incorporating additional triplets that capture concept relatedness.
Specifically, if two concepts $C_{i}$ and $C_{j}$ are in the same batch and $(C_{i},C_{j})\in\mathcal{R}$, we treat their definitions as additional positive pairs within the loss function, while using the definitions of all unrelated concepts as negatives. This setup implicitly introduces a ranking effect: definitions of the same concept are drawn closest together, as the corresponding loss triples include all other definitions - related or not - as negatives. In contrast, triples based on related concepts use only definitions of completely unrelated concepts as negatives, thereby encouraging related concepts to be embedded closer to one another than unrelated ones.
@@ 3.3 Pretraining Loss Combination
Our proposed loss is applied to the CLS token representation, allowing seamless integration with other pretraining losses that target the remaining token embeddings. This is especially interesting in light of recent models being trained exclusively with MLM loss warner2024modernbert, since the traditional next sentence prediction loss empirically did not lead to significant performance gains liu2019roberta. Consequently, our method presents a more sophisticated approach of infusing domain-relevant knowledge into the CLS token representation.
@@ 4 Ontology-Informed Pretraining
This section details the application and evaluation of our proposed method, using domain-specific ontologies for dataset creation. Our experiments focus on the scientific domain of invasion biology, a specialized subfield of biodiversity research that investigates non-native species, their introduction pathways, ecological impacts, and management strategies to mitigate their effects on ecosystems jeschke2018invasion.
@@ 4.1 Dataset Creation
Our approach involves constructing a domain-specific dataset consisting of concepts, definitions and concept relations in the target domain. To this end, we use two ontologies that address the target domain: the INBIO ontology alsayed_algergawy_2025_14864244, which captures concepts relevant to invasion biology, and the ENVO ontology buttigieg2013environment; buttigieg2016environment, which provides a structured representation of environmental and ecological concepts.
From these ontologies, we extract concept-definition pairs for all concepts that have a corresponding definition, as well as relational links between concepts. Additionally, we use a LLM, LLaMA-3-8B-Instruct grattafiori2024llama3herdmodels, to generate five additional definitions per concept, using the original ontology definition as context during generation to ensure that the new definitions accurately reflect the domain-specific meaning.
We compare our proposed pretraining approach to traditional MLM pretraining on sentences extracted from scientific abstracts. We leverage an existing index of paper titles in the field of invasion biology Mietchen2024ni and employ a web scraper to retrieve their abstracts, resulting in a final collection of 37,786 paper titles and abstracts.
@@ 4.2 Model Pretraining
In our experiments, we perform continual pretraining on a DeBERTa-base model he2021deberta by leveraging three different pretraining strategies:
@@
    Masked language modeling (MLM) pretraining with a masking probability of 0.25, applied to either abstract sentences, generated definitions, or a combined dataset of both.
@@
    Similarity (SIM) pretraining as described in Section 3, using our proposed similarity-based approach on the ontology-derived data.
@@
    Combined pretraining using MLM and SIM losses, done to investigate potential synergies between these approaches. We apply both strategies concurrently by performing two forward and backward passes - one for each loss function - for each parameter update.